The length of a rectangle is 6 inches more than its width. The perimeter is 24 inches. what is the length and width?
Answer by oscargut(2103)

Let length=L, width=W and perimeter=P
then
L=W+6
P=2(L+W)=24
P=2(W+6+W)=24
4W+12=24
4W=12
W=3
L=9